If Government of Canada deposits at a chartered bank (such as Scotia Bank) are transferred to a Government of Canada account at the Bank of Canada, the transfer will

a) enable chartered banks to create new money b) cause banks to decrease the quantity of money c) cause no change in the quantity of money.

29. Assume that Charisma is a closed economy with a marginal propensity to consume (MPC) of 0.8. If government expenditures increase by $100 billion, Charisma’s GDP will

a) increase by $100 billion b) decrease by $500 billion c) increase by $125 billion d) increase by $500 billion

30. Assume that Gamma is an open economy with a marginal propensity to consume (MPC) of 0.6 and a marginal propensity to import (MPM) of 0.1. If government expenditures increase by $100 billion, Gamma’s GDP will

a) increase by $100 billion b) decrease by $100 billion c) increase by $200 billion d) decrease by $200 billion
Answer #1

28. The right answer is a) enable chartered banks to create more money, which will be a function of money deposited by the Govt and the money multiplier

29. We can calculate increase in GDP = Increase in Govt exp / (1 - MPC) = 100 / (1 - 0.8) = 100 / 0.2 = $500B

So the right answer is d

30. We can calculate increase in GDP = Increase in Govt exp / (1 - MPC + MPM) = 100 / (1 - 0.6 + 0.1) = 100 / 0.5 = $200B

So the right answer is c